OpenStudy (anonymous):

find the y-intercept. y=-x^2+2x+4

OpenStudy (anonymous):

To find the y intercept, plug in 0 for x

OpenStudy (anonymous):

The y intercept of the graph is wherever you have an x value of 0. So, y=-x^2+2x+4 y=0+0+4

OpenStudy (anonymous):

Then find what y equals, and then put it like this (Value of y, 0)

OpenStudy (anonymous):

so (4,0)

OpenStudy (anonymous):

I am so sorry, It will be (0,4) I apologize

OpenStudy (anonymous):

lol ok thanks

OpenStudy (anonymous):

Here is a site http://www.purplemath.com/modules/intrcept.htm

OpenStudy (anonymous):

This site also gives review and helps finding x intercepts